# Fast unbiased sampling of networks with given expected degrees and strengths

### v3

@@ 2.6 Extension to other network types
Our sampling algorithm naturally extends to bipartite networks. We denote parameters for the one node set as $\alpha^{+},\beta^{+}$ and for the other node set as $\alpha^{-},\beta^{-}$. We then go through each node in the ++ node set and then go through each node in the - node set to sample the edge. The order of the nodes for the ++ node set and the - node set is sorted according to $\alpha^{+}_{i},\beta^{+}_{i}$ and $\alpha^{-}_{j},\beta^{-}_{j}$ respectively.
The sampling algorithm for bipartite networks can be used to generate directed networks as follows. Given a directed network, we create two nodes per node, one corresponding to the in nodes and the other corresponding to the out nodes, each placed in one part and the other (e.g., ++ part for in and - for out) in a bipartite network, with an edge from node $i$ to node $j$ in the directed network corresponds to an edge between out-node $i$ and in-node $j$ in the bipartite representation. This transformation allows us to apply the bipartite sampling algorithm to generate directed networks while preserving both in-degree and out-degree sequences. In the same vein, our algorithm can be applied to hypergraphs by representing them as incidence graphs.
@@ 3 Results
Figure 3: Results for the unweighted networks. A: CPU Time as a function of the number $M$ of edges. B: The joint probability distribution of the reconstructed and empirical degree sequences. We showed the results for three empirical networks as representative (Open Flight, ca-HepPh, and DBLP-cite). C: The density of edges within the group of the largest-degree nodes.
We analyzed 93 unweighted networks across domains and sizes, removing self-loops and multiedges (for network statistics, see Ref. [1]). We also used 10 weighted networks, with edge weights rounded to integers as required by Chung-Lu and UECM models [30]. To cover a wider range of network sizes, we converted the 93 unweighted networks to weighted networks by assigning one plus the number of common neighbors as edge weights, with one added to avoid zero weights. Parameters were inferred using NEMtropy [30] with quasi-newton optimization. Parameter inference time was excluded from sampling CPU measurements. We compared our algorithm against Brute Force (BF) sampling for MaxEnt models, MH for Chung-Lu (unweighted only), and stub-matching for Chung-Lu (weighted only). BF evaluates all node pairs using UBCM [30]. MH samples networks using Chung-Lu probabilities [19]. Stub-matching samples $C$ node pairs with probability proportional to node strength, where $C$ follows a Poisson distribution with mean equal to the sum of edge weights. Each algorithm ran 40 times per network, measuring CPU time excluding wait time.
@@ 3.2 Weighted networks
Figure 4: Results for the weighted networks. A: CPU Time as a function of the number $M$ of edges for weighted networks and synthetically-weighted networks. B: The joint probability distribution of the reconstructed and empirical degree sequences, and that of the strength sequences. We showed the results for three empirical networks as representative.
The proposed method samples faster while being as unbiased as the BF method for both originally weighted and synthetically weighted networks (Fig 4). The proposed method samples 10–1000 times faster than the BF and stub-matching methods for large networks, with the CPU time growing linearly with respect to the number of edges (Fig 4A). The BF and proposed methods preserved the expected degree and strength relatively well, with a notable deviation of the strength sequence for the Product Space SITC network. This is likely due to the inference error of the parameters of the UECM, not the sampling algorithms per se, as the BF method also shows a similar deviation. The stub-matching method based on the Chung-Lu model preserves the degree sequence well, but falls short for the degree sequence.
@@ 4 Discussion
We have developed a fast sampling algorithm for the MaxEnt models by adapting the MH algorithm originally designed for the Chung-Lu model. Our evaluation demonstrated computational speedups of 10-100 fold in CPU time for unweighted networks and 10—100 fold for weighted networks relative to the brute-force method, while preserving statistical accuracy.
Our approach echoes the combinatorial adaptation of the MH algorithm [28] that replaces the underlying edge probability derived from a combinatorial analysis of edge occurrences between two nodes. We did not compare against the combinatorial MH algorithm, as we focus on sampling efficiency rather than model differences. We note that both fast algorithms make it computationally feasible to systematically compare different configuration models on large, degree-heterogeneous networks where such comparisons were previously prohibitive.
Several limitations remain. First, the inference cost for MaxEnt models is still ${\cal O}(N^{2})$. Second, we did not identify exact theoretical computational complexity of our sampling algorithm as it depends on the parameter distribution, i.e., the algorithm becomes most efficient when $\beta$ values are homogeneous, approaching $\beta_{\min}\simeq\beta_{i}+\beta_{j}$ for all node pairs $i,j$, as this ensures tight upper bounds throughout the sampling process.
Despite these limitations, our proposed sampling algorithm improved the efficiency of unbiased configuration models. We hope that our sampling algorithm will bridge the gap between theoretical rigor and computational practicality.

### v4

@@ 2.6 Extension to other network types
Our sampling algorithm naturally extends to bipartite networks. We denote parameters for one node set as $\alpha^{+},\beta^{+}$ and for the other node set as $\alpha^{-},\beta^{-}$. We then go through each node in the ++ node set and then go through each node in the - node set to sample edges. The order of the nodes for the ++ node set and the - node set is sorted according to $\alpha^{+}_{i},\beta^{+}_{i}$ and $\alpha^{-}_{j},\beta^{-}_{j}$, respectively.
The sampling algorithm for bipartite networks can be used to generate directed networks as follows. Given a directed network, we create two nodes per node: one corresponding to the in nodes and the other corresponding to the out nodes, with one node type placed in one part of the bipartite network and the other node type in the other part. An edge from node $i$ to node $j$ in the directed network corresponds to an edge between out-node $i$ and in-node $j$ in the bipartite representation. This transformation allows us to apply the bipartite sampling algorithm to generate directed networks while preserving both in-degree and out-degree sequences. In the same vein, our algorithm can be applied to hypergraphs by representing them as incidence graphs.
@@ 3 Results
Figure 3: Results for the unweighted networks. A: CPU Time as a function of the number $M$ of edges. B: The joint probability distribution of the reconstructed and empirical degree sequences. We show the results for three empirical networks as representative (Open Flight, ca-HepPh, and DBLP-cite). C: The density of edges within the group of the largest-degree nodes.
We analyzed 93 unweighted networks across domains and sizes, removing self-loops and multi-edges (for network statistics, see Ref. [1]). We also used 10 weighted networks, with edge weights rounded to integers as required by the Chung-Lu and UECM models [28]. To cover a wider range of network sizes, we converted the 93 unweighted networks to weighted networks by assigning one plus the number of common neighbors as edge weights, with one added to avoid zero weights. Parameters were inferred using NEMtropy [28] with quasi-Newton optimization. Parameter inference time was excluded from sampling CPU measurements. We compared our algorithm against Brute Force (BF) sampling for MaxEnt models, MH for Chung-Lu (unweighted only), and strength-based sampling for Chung-Lu (weighted only). BF evaluates all node pairs using the UBCM [28]. MH samples networks using Chung-Lu probabilities [18]. Strength-based sampling first samples the total number of edges from a Poisson distribution with a mean equal to the sum of edge weights. Then, we attach each edge to a node pair with probability proportional to their node strengths. Each algorithm ran 40 times per network, measuring CPU time excluding wait time. We refer to our algorithm as FastMaxEnt.
@@ 3.2 Weighted networks
Figure 4: Results for the weighted networks. A: CPU Time as a function of $M$. B: The joint probability distribution of the reconstructed and empirical degree sequences, and that of the strength sequences. We show the results for three empirical networks as representative.
The FastMaxEnt method samples faster while being as unbiased as the BF method for both originally weighted and synthetically weighted networks (Fig. 4). The FastMaxEnt method samples 10–100 times faster than the BF and strength-based sampling methods for large networks, with CPU time growing linearly with respect to the number of edges (Fig. 4A). The BF and FastMaxEnt methods preserve the expected degree and strength relatively well, with a notable deviation of the strength sequence for the Product Space SITC network. This is likely due to inference error in the parameters of the UECM, not the sampling algorithms per se, as the BF method also shows a similar deviation. The strength-based sampling method based on the Chung-Lu model preserves the strength sequence well but falls short for the degree sequence.
@@ 4 Discussion
We have developed a fast sampling algorithm for MaxEnt models by adapting the MH algorithm originally designed for the Chung-Lu model. Our evaluation demonstrated computational speedups of 10–100 times in CPU time for unweighted networks and 10–100 times for weighted networks relative to the brute-force method while preserving statistical accuracy.
Our approach echoes the combinatorial adaptation of the MH algorithm [26] that replaces the underlying edge probability derived from combinatorial analysis of edge occurrences between two nodes. We did not compare against the combinatorial MH algorithm, as we focus on sampling efficiency rather than model differences. We note that both fast algorithms make it computationally feasible to systematically compare different configuration models on large, degree-heterogeneous networks where such comparisons were previously prohibitive.
Several limitations remain. First, the inference cost for MaxEnt models is still ${\cal O}(N^{2})$. Second, we did not identify the exact theoretical computational complexity of our sampling algorithm, as it depends on the parameter distribution. The algorithm becomes most efficient when $\beta$ values are homogeneous, approaching $\beta_{\min}\simeq\beta_{i}+\beta_{j}$ for all node pairs $i,j$, as this ensures tight upper bounds throughout the sampling process.
Despite these limitations, our proposed sampling algorithm improves the efficiency of unbiased configuration models. We hope that our sampling algorithm bridges the gap between theoretical rigor and computational practicality.